OpenStudy (anonymous):

simplify 3-x / |x-3|

OpenStudy (anonymous):

it will have ranged answer:

OpenStudy (anonymous):

Case I : \[\frac{ 3-x }{ |x-3| } = \frac{ 3-x }{ x-3 }\]

OpenStudy (anonymous):

for x>3 (3-x)/(x-3)=-1

OpenStudy (anonymous):

forx<3 (3-x)/(-(x-3))=1